Hello tutor,
Here is the question:
Find the domain of h(x)= -e^x
Description of transformation:
Equation(s) for the Horizontal asymptote(s):
y-intercept in (x, y) form:

Answer by stanbon(75887)

Find the domain of h(x)= -e^x
Domain: All Real Numbers
Description of transformation:
Start with y = e^x
Then y = e^-x is a reflection in the y-axis.
Then y = -e^-x is a reflection in the x-axis
-----------------------------------------------------
Equation(s) for the Horizontal asymptote(s):
HA is y = 0
-------------------------------------------
y-intercept in (x, y) form:
Let x = 0, then y = -e^0 = -1
In (x,y) form: (0,-1)